### v1

@@ 5.1 Finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$
Assume (H1) and (H2”). For every $k\geq 1$, $\widetilde{\mathscr{T}}^{n}_{k}\overset{d}{\longrightarrow}\mathcal{T}_{k}^{\ell,\sigma}$ in the sense of (6).
Proof 5.2.
The convergence in the sense of (6) is defined by the metric $d$ in (4). This metric is a function of the geometric structure of the tree, which is uniquely determined by the joint distribution of:
  • (1)
    the heights of the $k$ (ordered) leaves;
  • (2)
    the heights of all their associated branch points;
  • (3)
    the branching order of the tree.
Therefore, our goal is to show that the joint distribution of these components in $\widetilde{\mathscr{T}}^{n}_{k}$ converges to the corresponding joint distribution in $\mathcal{T}_{k}^{l}$. We break the proof into three parts.
(1) Convergence of leaf heights (”birth times”)
For any $\delta>0$, suppose $a_{n}\rightarrow a$ and $na_{n}$ are integers in $\delta n,h(q_{n})-\delta n]$. Then $n$ times the probability that the height of a uniform chosen vertex in $\mathscr{T}^{n}$ is $na_{n}$ is
$n\bigg(\sum_{s}q_{n}(s)\bigg)^{-1}q_{n}(na_{n})\longrightarrow\ell(a)/\int_{0}^{\infty}\ell(x)\,\mathrm{d}x.$ (80)
This resulting expression is precisely the probability density for the ”birth times” (i.e. heights) of the $k$ leaves in the limiting tree $\mathcal{T}_{k}^{l}$, as described in Theorem 2.20. By (H1), $n$ times the probability that the height of a uniform chosen vertex in $\mathscr{T}^{n}$ is $na_{n}$ in $0,\delta n]\cup[h(q_{n})-\delta n,h(q_{n})]$ goes to 0 as $\delta\rightarrow 0.$ Thus, the distribution of the $k$ leaf heights in $\widetilde{\mathscr{T}}_{k}^{n}$ converges to the distribution of leaf heights in $\mathcal{T}_{k}^{l}$.
(2) Convergence of branch point heights (”the coalescent process”)
Let $e_{n}(s)$ and $e(s)$ be the number of different branches at height $s$ of $\widetilde{\mathscr{T}}^{n}_{k}$ and $\mathcal{T}_{k}^{\ell}$ respectively. Denote by $a_{n,k}<\cdots<a_{n,2}<a_{n,1}$ the order statistics of the $k$ leaf heights of $\widetilde{\mathscr{T}}^{n}_{k}$. We first condition on the event that the $k$ leaf heights are in $\delta n,h(q_{n})-\delta n]$ for some $\delta$. Let us view $\widetilde{\mathscr{T}}^{n}_{k}$ as a coalescent process. First, we trace backward in time the lineage of the leaf at height $a_{n,1}$ until arriving at the height $a_{n,2}$. So $e_{n}(s)=1$ for $a_{n,2}<s\leq a_{n,,1}$. The ancestral at height $a_{n,2}$ of the first leaf is uniformly chosen among all the vertices at that height. So when $a_{n,3}<s\leq a_{n,2}$, $e_{n}(s)$ will behave like $X^{(2)}_{\lfloor ns\rfloor}(n,\lfloor na_{n,2}\rfloor)$. The coalescent process follows similarly below height $a_{n,3}$. By (1) and Proposition A.1, we can deduce that for every $\delta$ and $\delta<s<h(\ell)-\delta$,
$e_{n}(s)\rightarrow e(s).$ (81)
For the branch points near the boundary $h(q_{n})$, notice that the probability that branch points of $\mathscr{T}_{k}^{n}$ fall into $h(q_{n})-\delta,h(q_{n})]$ tends to 0 as $\delta\rightarrow 0$. For the branch points near the boundary 0, the difference of branch point heights of $\widetilde{\mathscr{T}}^{n}_{k}$ and $\mathcal{T}_{k}^{l}$ will tend to 0 as $\delta\rightarrow 0$. Thus, the joint distribution of the branch point heights in $\widetilde{\mathscr{T}}_{k}^{n}$ converges to that in $\mathcal{T}_{k}^{l}$.
(3) Convergence of branching order
$\mathscr{T}^{n}$ is defined by an exchangeable offspring distribution. This implies that the subtrees branching from any given parent are, in effect, placed in a uniformly random (lexicographic) order. This property directly corresponds to the construction of the limiting coalescent $\mathcal{T}_{k}^{l}$, where merging branches are randomly assigned a ”left” or ”right” position. Therefore, the order structure also converges.
@@
As is clear from the proof of Proposition 5.1, finite-dimensional convergence of the tree sequence is equivalent to convergence to Kingman’s coalescent in the sense of Proposition A.1.
@@
Assume (H1), (H2’) and (H2”). Then
$\big(n^{-1}C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})\big)_{t\geq 0}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(t)}^{\,\frac{4\ell}{\sigma^{2}}}\big)_{t\geq 0}\quad\text{in $D[0,\infty)$}.$ (82)
Proof $5.$
@@
$\big(n^{-1}C_{\lfloor 2(\#V(\mathscr{T}^{n})-1)t\rfloor}(\mathscr{T}^{n})\big)_{0\leq t\leq 1}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(I(\ell)t)}^{\,\frac{4\ell}{\sigma^{2}}}\big)_{0\leq t\leq 1}.$
Combining this with (2) gives (82).

### v2

@@ 5.1 Finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$
Assume (H1) and (H2b). Then for every $k\geq 1$,
[…]
Proof.
To begin with, we reduce the proof by controlling the heights of the sampled leaves. Fix $k\geq 1$. Let $\mathbf{a}_{1}\geq\mathbf{a}_{2}\geq\cdots\geq\mathbf{a}_{k}$ denote the heights of the leaves of the tree $\mathcal{T}^{\ell,\sigma}_{k}$ and $\mathbf{a}^{n}_{1}\geq\mathbf{a}^{n}_{2}\geq\cdots\geq\mathbf{a}^{n}_{k}$ be the order statistics of the set $\big\{\lvert\widetilde{V}^{n}_{1}\rvert,\ldots,\lvert\widetilde{V}^{n}_{k}\rvert\big\}$. For any $\delta>0$, define
$A^{n}_{\delta}$ $:=\big\{\mathbf{a}^{n}_{i}\in(\delta,h(\ell)-\delta)\ \forall\,i=1,\ldots,k\big\}\cap\big\{\mathbf{a}^{n}_{i}\neq\mathbf{a}^{n}_{j}\ \forall\,1\leq i\neq j\leq k\big\},$
$A_{\delta}$ $:=\big\{\mathbf{a}_{i}\in(\delta,h(\ell)-\delta)\ \forall\,i=1,\ldots,k\big\}\cap\big\{\mathbf{a}_{i}\neq\mathbf{a}_{j}\ \forall\,1\leq i\neq j\leq k\big\}.$
Then we have the following:
Lemma 5.2.
Assume (H1) and (H2b). Then for any $k\geq 1$ and $\eta>0$, there exists $\delta=\delta(k,\eta)>0$ such that
$\liminf_{n\rightarrow\infty}\mathbb{P}_{n}\left(A^{n}_{\delta}\right)>1-\eta,\quad\mathbb{P}(A_{\delta})>1-\eta.$ (5.2)
Proof.
The first bound follows directly from (H1) H1(i)H1(ii). The second bound follows from the piecewise Kingman representation in Proposition 2.14, together with the definition of continuous profiles in Section 1.1. ∎
Thanks to the lemma, it remains to prove for any $k\geq 1$ and $\eta>0$, there exists a coupling of $\Big(\big(\widetilde{\mathscr{T}}^{n}_{k},\{\widetilde{V}^{n}_{i}\}_{1\leq i\leq k}\big)\,\Big|\,A^{n}_{\delta}\Big)$, $n\geq 1$ and $\Big(\big(\mathcal{T}^{\ell,\sigma}_{k},\{L_{i}\}_{1\leq i\leq k}\big)\,\Big|\,A_{\delta}\Big)$ such that
$\mathbb{P}\left\{d\Big(\Big(\big(\widetilde{\mathscr{T}}^{n}_{k},\{\widetilde{V}^{n}_{i}\}_{1\leq i\leq k}\big)\,\Big|\,A^{n}_{\delta}\Big),\Big(\big(\widetilde{\mathscr{T}}^{n}_{k},\{\widetilde{V}^{n}_{i}\}_{1\leq i\leq k}\big)\,\Big|\,A^{n}_{\delta}\Big)\Big)\leq\eta\right\}\geq 1-2\eta,$ (5.3)
where $d$ denotes the distance in (2.1). Once (5.3) is verified, the desired convergence (5.1) follows from the following proposition:
Proposition 5.3.
Let $X$ and $(X_{n})_{n\geq 1}$ be random elements on a metric space $(S,\rho)$. If there exists a sequence $\Big\{\big(X^{k},(X^{k}_{n})_{n\geq 1}\big):k\geq 1\Big\}$ of couplings of $X$ and $(X_{n})_{n\geq 1}$ satisfying that for each $\varepsilon>0$,
$\lim_{k\rightarrow\infty}\limsup_{n\rightarrow\infty}{\mathbb{P}}\Big(\rho(X^{k},X^{k}_{n})>\varepsilon\Big)=0,$ (5.4)
then $X_{n}\overset{d}{\longrightarrow}X$.
Proof.
For each closed subset $F$ of $(S,\rho)$ and $\varepsilon>0$,
${\mathbb{P}}(X_{n}\in F)={\mathbb{P}}(X^{k}_{n}\in F)\leq{\mathbb{P}}\big(\rho(X^{k},X^{k}_{n})>\varepsilon\big)+{\mathbb{P}}(X\in F_{\varepsilon}),$
where $F_{\varepsilon}=\{x\in S:\rho(x,F)\leq\varepsilon\}$. Consequently,
$\limsup_{n\rightarrow\infty}{\mathbb{P}}(X_{n}\in F)\leq\lim_{k\rightarrow\infty}\limsup_{n\rightarrow\infty}{\mathbb{P}}\Big(\rho(X^{k},X^{k}_{n})>\varepsilon\Big)+{\mathbb{P}}(X\in F_{\varepsilon})={\mathbb{P}}(X\in F_{\varepsilon}).$
Letting $\varepsilon\downarrow 0$ yields
$\limsup_{n\rightarrow\infty}{\mathbb{P}}(X_{n}\in F)\leq{\mathbb{P}}(X\in F),$
which is equivalent to $X_{n}\overset{d}{\longrightarrow}X$. ∎
For the proof of (5.3), we introduce the following notation. For $s\in[\delta,h(\ell)-\delta]$, let
$Y_{s}=Y_{s}^{(k)}$ $:=\#\,\text{ancestors of $L_{1},\ldots,L_{k}$ in $\mathcal{T}^{\ell,\sigma}_{k}$ at height $s$},$
$Y^{n}_{s}=Y_{s}^{n,(k)}$ $:=\#\,\text{ancestors of $\widetilde{V}^{n}_{1},\ldots,\widetilde{V}^{n}_{k}$ in $\widetilde{\mathscr{T}}^{n}_{k}$ at height $\lceil ns\rceil/n$}.$
The proof is accomplished by a chain of lemmas.
Lemma 5.4.
For any $k\geq 1$ and $\delta>0$,
$\big((\mathbf{a}^{n}_{1},\ldots,\mathbf{a}^{n}_{k})\,\big|\,A^{n}_{\delta}\big)\overset{d}{\longrightarrow}\big((\mathbf{a}_{1},\ldots,\mathbf{a}_{k})\,\big|\,A_{\delta}\big).$
Proof.
This follows directly from (H1), (1.1) and the first step of the equivalence construction of $\mathcal{T}^{\ell}_{k}$ presented after Proposition 2.14. ∎
Lemma 5.5.
Fix $k\geq 1$ and $\delta>0$. Let $(a_{1}^{n},\ldots,a_{k}^{n})\in\big((\delta,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}\big)^{k}$ for $n\geq 1$, and $(a_{1},\ldots,a_{k})\in(\delta,h(\ell)-\delta)^{k}$ satisfy
$(a_{1}^{n},\ldots,a_{k}^{n})\longrightarrow(a_{1},\ldots,a_{k})\quad\text{and}\quad a_{1}>\cdots>a_{k}.$
Then
$\begin{aligned} \Big(\big(Y^{n}(s):s\in[\delta,h(\ell)-\delta]\big)\,\big|\,(\mathbf{a}^{n}_{1},\ldots,\mathbf{a}^{n}_{k})=(a^{n}_{1},\ldots,a^{n}_{k})\Big)\\ \overset{d}{\longrightarrow}\Big(\big(Y(s):s\in[\delta,h(\ell)-\delta]\big)\,\big|\,(\mathbf{a}_{1},\ldots,\mathbf{a}_{k})=(a_{1},\ldots,a_{k})\Big).\end{aligned}$ (5.5)
Lemma 5.6.
Let $\delta>0$, and let
$y^{n}=\big(y^{n}(s):s\in[\delta,h(\ell)-\delta]\big),\qquad y=\big(y(s):s\in[\delta,h(\ell)-\delta]\big)$
be the deterministic paths in the value spaces of $(Y^{n}_{s}\,|\,A^{n}_{\delta})$ and $(Y_{s}\,|\,A_{\delta})$, respectively, such that
[…]
Then there exists a coupling of $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$, $n\geq 1$, and $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$ such that a.s.,
$\limsup_{n\rightarrow\infty}d\Big\{\Big(\big(\widetilde{\mathscr{T}}^{n}_{k},\{\widetilde{V}^{n}_{i}\}_{1\leq i\leq k})\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)\Big),\Big(\big(\mathcal{T}^{\ell,\sigma}_{k},\{L_{i}\}_{1\leq i\leq k})\,\big|\,A_{\delta},\;Y=y\big)\Big)\Big\}\leq 2k\delta.$
Proof of (5.3) assuming Lemmas 5.5 and 5.6.
It follows from Lemmas 5.4 and 5.5 that
$(Y^{n}\mid A^{n}_{\delta})\;\overset{d}{\longrightarrow}\;(Y\mid A_{\delta})\qquad\text{in }D[\delta,h(\ell)-\delta].$
Moreover, by the piecewise Kingman representation of $Y$ in Proposition 2.14, we have a.s., $(Y\,|\,A_{\delta})$ has jumps of size at most 11. Thus, further combining Lemma 5.6 with sufficiently small $\delta>0$ yields (5.3). ∎
We now turn to the proof of Lemmas 5.5 and 5.6.
Proof of Lemma 5.5.
We begin by introducing a preparatory lemma, for which we set up some notation.
Let $(b_{1}^{n},\ldots,b_{m}^{n})\in\big((\delta,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}\big)^{m}$ for $n\geq 1$, and $(b_{1},\ldots,b_{m})\in(\delta,h(\ell)-\delta)^{m}$ satisfy
$(b_{1}^{n},\ldots,b_{m}^{n})\longrightarrow(b_{1},\ldots,b_{m}),\qquad b^{n}_{1}>\cdots>b^{n}_{m},\qquad b_{1}>\cdots>b_{m},$
$\{b^{n}_{1},\ldots,b^{n}_{m}\}\cap\{a^{n}_{1},\ldots,a^{n}_{k}\}=\emptyset,\quad\{b_{1},\ldots,b_{m}\}\cap\{a_{1},\ldots,a_{k}\}=\emptyset.$
Define the deterministic paths $y^{n}=\big(y^{n}(s):s\in[\delta,h(\ell)-\delta]\big)$ (resp. $y=\big(y(s):s\in[\delta,h(\ell)-\delta]\big)$) as follows: viewed backwards from $h(\ell)-\delta$ down to $\delta$, the path starts at $y^{n}(h(\ell)-\delta)=0$ (resp. $y(h(\ell)-\delta)=0$), has upward jumps of size +1+1 at times $a_{1}^{n},\ldots,a_{k}^{n}$ (resp. $a_{1},\ldots,a_{k}$), and downward jumps of size 1-1 at times $b_{1}^{n},\ldots,b_{m}^{n}$ (resp. $b_{1},\ldots,b_{m}$). (If a time $a_{i}^{n}$ coincides with some $b_{j}^{n}$ (resp. $a_{i}$ with $b_{j}$), then the two jumps cancel and the path has no jump at that time.)
For any $s\in(\delta,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}$, define the event
$B_{s}=B_{s}(n,\delta):=\left\{Y^{n}|_{[s,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}}=y^{n}|_{[s,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}}\right\}\cap A^{n}_{\delta}.$ (5.6)
We now state the lemma (its proof is postponed):
Lemma 5.7.
For any $\delta>0$, $k\geq 1$, and sequences $(a_{i}^{n})$, $(a_{i})$, $(b_{i}^{n})$, $(b_{i})$ satisfying the above assumptions, we have uniformly for all $s\in(\delta,h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}$,
$\mathbb{P}_{n}\left(Y^{n}(s-\tfrac{1}{n})=y^{n}(s-\tfrac{1}{n})\,|\,B_{s}\right)=\,\begin{cases}1+o_{n}(1),&\quad\text{if }s-\tfrac{1}{n}\in\{a^{n}_{1},\ldots,a^{n}_{k}\},\\ \dfrac{\binom{y^{n}(s)}{2}\sigma_{n}(s)^{2}}{q_{n}(s)}\big(1+o_{n}(1)\big),&\quad\text{if }s-\tfrac{1}{n}\in\{b^{n}_{1},\ldots,b^{n}_{m}\},\\ 1-\dfrac{\binom{y^{n}(s)}{2}\sigma_{n}(s)^{2}}{q_{n}(s)}\big(1+o_{n}(1)\big),&\quad\text{otherwise}.\end{cases}$
Let us use the lemma to prove Lemma 5.5. Let
$p_{n}:=\mathbb{P}_{n}(Y^{n}=y^{n}\,|\,A^{n}_{\delta}),$
and let $p$ denote the density of the probability that the conditioned process $(Y\,|\,A_{\delta})$ follows the same jump pattern as $y$, with upward jump times perturbed by $\mathrm{d}a_{1},\ldots,\mathrm{d}a_{k}$.
It follows from Proposition 2.14 that a.s.,
$\{b_{1},\ldots,b_{m}\}\cap\{a_{1},\ldots,a_{k}\}=\emptyset.$
Thus by [4, Theorem 3.3], it is enough to show the following local limit estimate: for any $\delta>0$, $k\geq 1$, and sequences $(a_{i}^{n})$, $(a_{i})$, $(b_{i}^{n})$, $(b_{i})$ satisfying the above assumptions,
$n^{k}\,p_{n}\;\longrightarrow\;p.$ (5.7)
To this end, we first note that
$\{Y^{n}=y^{n}\}=\Big\{\,Y^{n}\big|_{(\delta,h(\ell)-\delta)\cap\tfrac{1}{n}\mathbb{Z}}=y^{n}\big|_{(\delta,h(\ell)-\delta)\cap\tfrac{1}{n}\mathbb{Z}}\,\Big\},$
by the definition of $Y^{n}$. Thus, it follows from Lemma 5.7 and (H1) that
$\begin{aligned} n^{k}p_{n}&=n^{k}\prod_{s\in(\delta+\frac{1}{n},h(\ell)-\delta)\cap\frac{1}{n}\mathbb{Z}}\mathbb{P}_{n}\left(Y^{n}(s-\tfrac{1}{n})=y^{n}(s-\tfrac{1}{n})\,|\,B_{s}\right)\\ &\underset{n\rightarrow\infty}{\longrightarrow}\,\exp\left[-\int_{\delta}^{h(\ell)-\delta}\binom{y(s)}{2}\frac{\sigma(s)^{2}}{\ell(s)}\,\mathrm{d}s\right]\prod_{i=1}^{k}\binom{y(b_{i})}{2}\dfrac{\sigma(b_{i})^{2}}{\ell(b_{i})}.\end{aligned}$ (5.8)
On the other hand, by Proposition 2.14, $(Y\,|\,(\mathbf{a}_{1},\ldots,\mathbf{a}_{k})=(a_{1},\ldots,a_{k}))$ is distributed as the piecewise Kingman’s coalescent $Z_{s}^{(k)}(a_{1},\ldots,a_{k})$. A direct inspection of its transition mechanism shows that the RHS of (5.8) is exactly the path density $p$. This proves (5.7). ∎
Proof of Lemma 5.6.
The proof also requires a preliminary lemma.
Assume the same setting as in Lemma 5.7. Let $\mathbf{W}^{n}_{i}$ (resp. $\mathbf{W}_{i}$), $1\leq i\leq m$, be the leaf of $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$ (resp. $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$) at height $a^{n}_{i}$ (resp. $a_{i}$), and $\mathbf{B}^{n}_{i}$ (resp. $\mathbf{B}_{i}$), $1\leq i\leq k$, be the branching point of $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$ (resp. $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$) at height $b^{n}_{i}$ (resp. $b_{i}$).
Lemma 5.8.
Assume the same setting as in Lemma 5.7. Then the tree order of the leaves $(\mathbf{W}^{n}_{i})$ and branching points $(\mathbf{B}_{i}^{n})$ of $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$ are uniform over all possible orders. The analogous result also holds for $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$.
We postpone the proof of the lemma. We now work under the assumptions of Lemma 5.6. Observe that the convergence $y^{n}\overset{d}{\longrightarrow}y$ in the statement of Lemma 5.6 is equivalent to the convergence of the corresponding upward and downward jump times. Thus Lemma 5.8 implies that for sufficiently large $n$, there exists a coupling of $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$ and $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$ such that the tree order of their leaves and branching points agree. That is,
$\mathbf{W}^{n}_{i}\prec\mathbf{W}_{j}\text{ iff }\mathbf{W}_{i}\prec\mathbf{W}_{j},\qquad\mathbf{B}_{i}^{n}\prec\mathbf{B}_{j}^{n}\text{ iff }\mathbf{B}_{i}\prec\mathbf{B}_{j},\qquad\mathbf{W}^{n}_{i}\prec\mathbf{B}_{j}^{n}\text{ iff }\mathbf{W}_{i}\prec\mathbf{B}_{j}.$
Consequently, let $\mathbf{W}^{n}_{(1)},\ldots,\mathbf{W}^{n}_{(m)}$ (resp. $\mathbf{W}_{(1)},\ldots,\mathbf{W}_{(m)}$) denote the order statistic of $\mathbf{W}^{n}_{1},\ldots,\mathbf{W}^{n}_{m}$ (resp. $\mathbf{W}_{1},\ldots,\mathbf{W}_{m}$). Then for any $1\leq i\leq m-1$,
  1. 1.
    the branching point of $\mathbf{W}^{n}_{(i)}$ and $\mathbf{W}^{n}_{(i+1)}$ belongs to $\{\mathbf{B}_{j}^{n}\}$ (i.e. the height of branching points is in $\delta,h(\ell)-\delta]$) iff the branching point of $\mathbf{W}_{(i)}$ and $\mathbf{W}_{(i+1)}$ belongs to $\{\mathbf{B}_{j}\}$;
  2. 2.
    $\mathbf{B}_{j}^{n}$ is the branching point of $\mathbf{W}^{n}_{(i)}$ and $\mathbf{W}^{n}_{(i+1)}$ iff $\mathbf{B}_{j}$ is the branching point of $\mathbf{W}_{(i)}$ and $\mathbf{W}_{(i+1)}$.
Now recall the distance defined in (2.1). For $1\leq i\leq m-1$, writing $\mathbf{B}^{n}_{(i)}$ (resp. $\mathbf{B}_{(i)}$) for the branching point of $\mathbf{W}^{n}_{(i)}$ and $\mathbf{W}^{n}_{(i+1)}$ (resp. $\mathbf{W}_{(i)}$ and $\mathbf{W}_{(i+1)}$), we distinguish two cases:
  1. 1.
    if both $\mathbf{B}^{n}_{(i)}$ and $\mathbf{B}_{(i)}$ have heights in $(\delta,h(\ell)-\delta)$, then for some $j$, $\mathbf{B}^{n}_{(i)}=\mathbf{B}_{j}^{n}$ and $\mathbf{B}_{(i)}=\mathbf{B}_{j}$. Using the convergence of the heights of the leaves and branching points, we obtain a.s.
    $d(\mathbf{W}^{n}_{(i)},\mathbf{B}^{n}_{(i)})\longrightarrow d(\mathbf{W}_{(i)},\mathbf{B}_{(i)}),\qquad d(\mathbf{W}^{n}_{(i+1)},\mathbf{B}^{n}_{(i)})\longrightarrow d(\mathbf{W}_{(i+1)},\mathbf{B}_{(i)}),$
    where $d$ denotes the tree distance in the corresponding space (either on $\big(\widetilde{\mathscr{T}}^{n}_{k}\mid A^{n}_{\delta},\,Y^{n}=y^{n}\big)$ or on $\big(\mathcal{T}^{\ell,\sigma}_{k}\mid A_{\delta},\,Y=y\big)$, respectively);
  2. 2.
    otherwise, both $\mathbf{B}^{n}_{(i)}$ and $\mathbf{B}_{(i)}$ lie in $0,\delta]$, and hence
    $\limsup_{n\rightarrow\infty}\left\lvert d(\mathbf{W}^{n}_{(i)},\mathbf{B}^{n}_{(i)})-d(\mathbf{W}_{(i)},\mathbf{B}_{(i)})\right\rvert\leq 2\delta,\qquad\limsup_{n\rightarrow\infty}\left\lvert d(\mathbf{W}^{n}_{(i+1)},\mathbf{B}^{n}_{(i)})-d(\mathbf{W}_{(i)},\mathbf{B}_{(i)})\right\rvert\leq 2\delta.$
It is also clear that
$d(\mathbf{W}^{n}_{(1)},\text{root})\longrightarrow d(\mathbf{W}_{(1)},\text{root}),\quad d(\mathbf{W}^{n}_{(m)},\text{root})\longrightarrow d(\mathbf{W}_{(m)},\text{root}).$
Putting everything together, we obtain Lemma 5.6 under the above coupling. ∎
Finally, we verify Lemmas 5.7 and 5.8. Recall the definition of $B_{s}$ in (5.6). By recursively applying Proposition 3.8 downward from height $h(\ell)-\delta$, we obtain that for every $s\in(\delta,h(\ell)-\delta)\cap\tfrac{1}{n}\mathbb{Z}$, conditionally on $B_{s}$, the $y^{n}(s)$ ancestors of $\widetilde{V}^{n}_{1},\ldots,\widetilde{V}^{n}_{k}$ at height $s$ are distributed as a uniformly chosen $y^{n}(s)$-tuple of vertices at that height.
Proof of Lemma 5.7.
Using Lemma 4.12, we have for all $s\in(\delta,h(\ell)-\delta)\cap\tfrac{1}{n}\mathbb{Z}$,
$\mathbb{P}_{n}\!\left(Y^{n}(s-\tfrac{1}{n})=y^{n}(s-\tfrac{1}{n})\,\middle|\,B_{s}\right)$
$=\;$ $\begin{cases}\displaystyle\Big(1-\frac{y^{n}(s)}{q^{n}(ns)}\Big)\mathbb{P}_{n}\!\left(\#\{\mathfrak{p}(u^{ns}_{1}),\ldots,\mathfrak{p}(u^{ns}_{y^{n}(s)})\}=y^{n}(s)\right),&\text{if }s-\tfrac{1}{n}\in\{a^{n}_{1},\ldots,a^{n}_{k}\},\\ \displaystyle\mathbb{P}_{n}\!\left(\#\{\mathfrak{p}(u^{ns}_{1}),\ldots,\mathfrak{p}(u^{ns}_{y^{n}(s)})\}=y^{n}(s)-1\right),&\text{if }s-\tfrac{1}{n}\in\{b^{n}_{1},\ldots,b^{n}_{m}\},\\ \displaystyle\mathbb{P}_{n}\!\left(\#\{\mathfrak{p}(u^{ns}_{1}),\ldots,\mathfrak{p}(u^{ns}_{y^{n}(s)})\}=y^{n}(s)\right),&\text{otherwise}.\end{cases}$
Here, the factor $1-\tfrac{y^{n}(s)}{q^{n}(ns)}$ is the probability that the newly created leaf at height $s-\tfrac{1}{n}$ avoids the $y^{n}(s)$ ancestral lineages above height $s$. Combining the above with (3.14) and (3.15) yields the desired result. ∎
Proof of Lemma 5.8.
The statement for $\big(\mathcal{T}^{\ell,\sigma}_{k}\,\big|\,A_{\delta},\;Y=y\big)$ is a consequence of the final step in the piecewise Kingman’s construction described in Proposition 2.14.
We now treat $\big(\widetilde{\mathscr{T}}^{n}_{k}\,\big|\,A^{n}_{\delta},\;Y^{n}=y^{n}\big)$. List the $y^{n}(s)$ vertices at height $s$ in tree order as
$z^{ns}_{i_{1}}\prec z^{ns}_{i_{2}}\prec\cdots\prec z^{ns}_{i_{y^{n}(s)}}.$
In the application of Lemma 4.12 that appeared in the proof of Lemma 5.7, we additionally keep track of the probability that:
  1. 1.
    when $s\in\{a^{n}_{1},\ldots,a^{n}_{k}\}$, the newly created leaf is exactly $z^{ns}_{i_{j}}$;
  2. 2.
    when $s\in\tfrac{1}{n}+\{b^{n}_{1},\ldots,b^{n}_{m}\}$, the vertices $z^{ns}_{i_{j}}$ and $z^{ns}_{i_{j+1}}$ are the pair that coalesces at height $s$.
In each case, one checks that the index $j$ is uniformly distributed over all its possible values (that is, over $\{1,\ldots,y^{n}(s)\}$ in the first case and $\{1,\ldots,y^{n}(s)-1\}$ in the second).
Moreover, a straightforward combinatorial check shows that the possible choices of $j$ at each jump time $s$ are in bijection with all possible tree orders of the leaves and branch points. Hence all such orders occur with equal probability. ∎
@@
Under (H1), the following three statements are in fact equivalent:
  1. (i)
    the finite-dimensional convergence in Proposition 5.1;
  2. (ii)
    convergence of the ancestral process to Kingman’s coalescent;
  3. (iii)
    the third-moment condition (H2b).
Indeed, we have already shown that (H2b) implies the finite-dimensional convergence, and hence (iii) implies both (i) and (ii). Conversely, if (H2b) fails, then along any subsequence $(n_{k})$ for which the limit in (H2b) is strictly positive, the probability of a triple–merger event remains bounded away from 0. This prevents the ancestral process from converging to Kingman’s coalescent and therefore rules out the finite-dimensional convergence.
@@
Assume (H1)–(H2). Then
$\big(n^{-1}C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})\big)_{t\geq 0}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(t)}^{\,\ell^{\sigma}}\big)_{t\geq 0}\quad\text{in $D[0,\infty)$}.$ (5.9)
Proof.
@@
$\big(n^{-1}C_{\lfloor 2(\#V(\mathscr{T}^{n})-1)t\rfloor}(\mathscr{T}^{n})\big)_{0\leq t\leq 1}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(I(\ell)t)}^{\,\ell^{\sigma}}\big)_{0\leq t\leq 1}.$
Combining this with (1.1) gives (5.9). ∎